OpenStudy (anonymous):

find critical points of (x+1)/(x^(2)+1)

OpenStudy (anonymous):

i got the derivative as [(x^(2)+1)-(x+1)(2x)]/[(x^(2)+1]^2

OpenStudy (anonymous):

Other than starting a new topic because you didn't like my answer... did you do that again? And would you like me to do it step by step?

OpenStudy (anonymous):

i changed the derivative because you said it wasn't right. you should probably check before accusing someone.

OpenStudy (anonymous):

You started a new topic. It made me feel sad. Give me a second. This has a lot of little stuffs.

OpenStudy (anonymous):

the derivative is right.

OpenStudy (anonymous):

what should i do next?

OpenStudy (anonymous):

I've totally messed this question up. I'm sorry. I'm leaving it to someone elses capable hands.

OpenStudy (anonymous):

find the roots of each side of the fraction. the roots are your critical numbers.

OpenStudy (anonymous):

you mean solve for x? like (x^(2)+1)-(x+1)(2x)=0?

OpenStudy (anonymous):

i tried that but got with x^2 = 2x-1

OpenStudy (anonymous):

* got stuck

OpenStudy (anonymous):

Oh. I can do that part. x^2=2x-1 x^2-2x+1 (x-1)(x-1) x=1

OpenStudy (anonymous):

that looks like a quadratic equation, where quadratic formula comes handy.

OpenStudy (anonymous):

so 1 is the critical point? the only critical point?

OpenStudy (anonymous):

the numerator is equivalent to this \[\huge x^2+1=(x+1)(2x)\]

OpenStudy (anonymous):

did you mean the denominator?

OpenStudy (anonymous):

the numerator. \[\large (x^2+1)-(x+1)(2x) = 0\]

OpenStudy (anonymous):

yes. and then if you it for x you get (x-1)(x-1) using quadratics. so 1 is a critical point?

OpenStudy (anonymous):

i think the quadratic equation is \[\huge -x^2 \large - 2x + 1=0\]

OpenStudy (anonymous):

yup. you are right. i missed a negative. now i get -2.414 and .4142

OpenStudy (anonymous):

that's \[\large -1 \pm \sqrt{2} \] all right.

OpenStudy (anonymous):

all right

OpenStudy (anonymous):

so would those be the critical points?

OpenStudy (anonymous):

just the two of them. the denominator yield no critical number.

OpenStudy (anonymous):

ok. thank you

OpenStudy (anonymous):

yw; and thank your for the medal.

OpenStudy (anonymous):

@r0n4ld i was just checking the answers and i plugged -1 and sqrt(2) in the function. i got 0 if i plugged in -1 but sqrt(2)

OpenStudy (anonymous):

*but not sqrt(2)

OpenStudy (anonymous):

your roots in decimal form is the approximation to the exact roots, that is \[ \large -2.414 \approx -1 - \sqrt{2}\] and \[\large 0.4142 \approx -1 + \sqrt{2} \]

OpenStudy (anonymous):

so you mean that if i plug in -2.414 and .04142 then it'll be 0 just to confirm.

OpenStudy (anonymous):

yes, the answer is zero in \[\huge f \prime(x)\] NOT \[\huge f(x)\]

OpenStudy (anonymous):

oh. ok. i'll check it again then